OpenStudy (anonymous):

write the standard equation of a cirlce with the center -2,-4 and radius: 10

OpenStudy (anonymous):

(x+2)^2 +(y+4)^2 = 10^2

OpenStudy (anonymous):

okay what about the x and y intercepts for y= -I(abs.value) x+5I

OpenStudy (anonymous):

for x- intercept put y=0 and for y-intercept put x=0 in y= -I(abs.value) x+5I

OpenStudy (anonymous):

so would the equation just be y= x+5 since theres an absolute value mark on each side

OpenStudy (anonymous):

dont consider the absolute

OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line through the given point and parallel to the given line. point : (3/5, 1/2) & line is 2x+5y=0

OpenStudy (anonymous):

consider this equation first 2x+5y=0 can you write the slope intercept form of this equation??

OpenStudy (anonymous):

isn't it y=-2/5x+0 ?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so what is the slope here???

OpenStudy (anonymous):

-2/5, now what

OpenStudy (anonymous):

ok you got the slope = -2/5 now two lines have the same slope when they are parallel thus the slope for the line you want to find is same as -2/5 now you have the slope = -2/5 and one point (3/5, 1/2) use point slope form now to get the equation

OpenStudy (anonymous):

isn't the point slope form y-y1=m(x-x1) ???

OpenStudy (anonymous):

yesssssss

OpenStudy (anonymous):

k.one minute don't go anywhere I have more questions after this!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

right now im at y- 1/2= -2/5x + 6/25. I have to retrice1/2 to 6/25 but idk how to add those bc what could even be the common denominator to add it?

OpenStudy (anonymous):

retrace mean ADD idk what that word was

OpenStudy (anonymous):

ok its right now leave y alone what you get??

OpenStudy (anonymous):

when I leave y alone I get : y= -2/5 + (6/25+1/2) how do I add those last two

OpenStudy (anonymous):

by LCM

OpenStudy (anonymous):

i know i have to find a common denominator but with 2 and 25 that's not possible

OpenStudy (anonymous):

take the LCM of 25 and 2 first

OpenStudy (anonymous):

thatd be 1 right?

OpenStudy (anonymous):

bc 2 doesn't go into 25

OpenStudy (anonymous):

no its 50

OpenStudy (anonymous):

so would i multiply the numerator by 50? im confused at how this will be set up with a lcm of 50

OpenStudy (anonymous):

ok i will draw the this step for you

OpenStudy (anonymous):